# Inverse scattering for $N$-body time-decaying harmonic oscillators

### v1

@@ 1 Introduction
We now give the assumptions for each potential function $V_{jk}$ in (1.1) as a perturbation of $H_{0}(t)$. Let the condition for $\lambda$ be
@@
with $\rho>1/(1-\lambda)$, where $\langle\cdot\rangle=\sqrt{1+|\cdot|}$ and $A\lesssim B$ means that there exists a constant $C>0$ such that $A\leqslant CB$. The singular part $V_{jk}^{\rm sing}\in L^{q}(\mathbb{R}^{d})$ is compactly supported on $\mathscr{X}$, where the Lebesgue exponent $q$ satisfies
@@
acting on $L^{2}(\mathscr{X})$. Here we note $r_{k}-r_{j}\in\mathscr{X}$ for any $1\leqslant j<k\leqslant N$. Under these definitions, the perturbed full Hamiltonian is decomposed into
@@
Many researchers have studied scattering theory under the $N$-body quantum systems, in particular, the the existence and asymptotic completeness of the wave operators, has been studied by a lot of researchers. Regarding the standard $N$-body Schrödinger operators, see [5, 13, 14, 18, 46, 47, 48]; the detailed history and proofs are also provided in [15, 31]. For $N$-body Schrödinger operators in the external electric fields, see [1, 3, 8, 9, 20]. For the external magnetic fields, see [2, 17]. However, there are no results regarding scattering theory for $N$-body time-decaying harmonic oscillators.
@@
Regarding scattering theory for the two-body time-decaying harmonic oscillator, [28] proved the existence of the wave operators under the assumption 1.1 only for the bounded part $V^{\rm bdd}$ and non-existence for $\rho\leqslant 1/(1-\lambda)$, and clarified that the condition $\rho>1/(1-\lambda)$ is short-range. We can intuitively understand the threshold $1(1-\lambda)$ from the classical motion of the particle $x(t)=c_{1}t^{1-\lambda}+c_{2}t^{\lambda}$ for $t\geqslant T$, which satisfies the Newton equation $({\rm d}^{2}/{\rm d}t^{2})x(t)=-k(t)x(t)$ because $\lambda$ is one of the roots of the quadratic equation $\lambda^{2}-\lambda+\sigma=0$. The critical case $\sigma=1/4$ was studied by [29]. For the inverse square potential, [30] proved the asymptotic completeness of the wave operators. Moreover, [30] proved the Strichartz estimates and applied them to the initial-value problems for the non-linear Schrödinger equations. The non-linear Schrödinger equations and Strichartz estimates with time-decaying harmonic oscillators have also been studied by [33, 34, 35, 36, 37, 38, 39, 40].
@@ 2 Existence of Wave Operators
$\|F(\sqrt{\mu_{jk}}|r_{k}-r_{j}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))e^{-{\rm i}t^{1-2\lambda}D^{2}/(2(1-2\lambda))}\phi\|\lesssim_{\nu}t^{-\nu(1-2\lambda)}\|\langle r_{k}-r_{j}\rangle^{\nu}\phi\|$ (2.11)
holds for any $1\leqslant j<k\leqslant N$, where $\mu_{jk}=m_{j}m_{k}/(m_{j}+m_{k})$, $\nu\in\mathbb{N}$, and $t\geqslant T$, and $\lesssim_{\nu}$ means that the constant depends on $\nu$.
@@
where the Jacobi coordinates are reconstructed from $y_{1}=r_{k}-r_{j}$ for $(j,k)\not=(1,2)$. Because $|\xi_{1}|=|L_{jlk}\xi|\geqslant\epsilon_{0}|\xi|\geqslant\epsilon^{2}$ holds on $\operatorname{supp}\mathscr{F}_{d(N-1)}\phi$, it suffices to prove that
$\|F(|x_{1}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))e^{-{\rm i}t^{1-2\lambda}D_{1}^{2}/(2(1-2\lambda))}\psi\|_{L^{2}(\mathbb{R}^{d})}\lesssim_{\nu}t^{-\nu(1-2\lambda)}\|\langle x_{1}\rangle^{\nu}\psi\|_{L^{2}(\mathbb{R}^{d})}$ (2.13)
@@
and integrating by parts, we have (2.13). ∎
@@
$\lesssim t^{-\nu(1-2\lambda)}\|\langle x\rangle^{\nu}\phi\|$ (2.20)
@@
$\chi(x)=\begin{cases}\ 1&\mbox{if}\quad|x|\geqslant 1\\ \ 0&\mbox{if}\quad|x|\leqslant 1/2.\end{cases}$ (2.21)
@@
We can use Proposition 2.2 again for the first and second terms on the right side of (2.3). By (2.18), (2.19), (2.20), and (2.24), we therefore have
$\|({\rm d}/{\rm d}t)U_{\lambda}(t,T)^{*}e^{-{\rm i}t^{1-2\lambda}D^{2}/(2(1-2\lambda))}\phi\|\lesssim t^{-\rho(1-\lambda)}+t^{-\nu(1-2\lambda)}+t^{-2+2\lambda}$ (2.25)
for $t\gg 1$. We note that $\rho(1-\lambda)>1$, $2-2\lambda>1$, and $\nu(1-2\lambda)>1$ for $\nu\gg 1$. By the Cook-Kuroda method ([45, Theorem XI.4]), this completes the proof. ∎
@@ 3 Inverse Scattering
For $v\in\mathbb{R}^{d}$, the normalization is $\hat{v}=v/|v|$. Let $\hat{v},e_{3},\ldots,e_{N}\in\mathbb{R}^{d}$ be unit vectors pointing in mutually different directions. We define $v_{jk}=v_{k}-v_{j}$ for $1\leqslant j<k\leqslant N$ where $v_{1}=-v/2$, $v_{2}=v/2$, and $v_{j}=|v|^{2}e_{j}$ for $3\leqslant j\leqslant N$. Let $\Phi_{0}\in\mathscr{S}(\mathbb{R}^{d(N-1)})$ be $\Phi_{0}=\phi_{1}\otimes\phi_{2}$ such that $\mathscr{F}_{d}\phi_{1}\in C_{0}^{\infty}(\mathbb{R}^{d})$ and $\mathscr{F}_{d(N-1)}\phi_{2}\in C_{0}^{\infty}(\mathbb{R}^{d(N-2)})$ with $\|\phi_{2}\|_{L^{2}(\mathbb{R}^{d(N-2)})}=1$. In addition, we assume that $\operatorname{supp}\mathscr{F}_{d}\phi_{1}\subset\{\xi_{12}\in\mathbb{R}^{d}\bigm||\xi_{12}|<\eta_{12}\}$ for some $\eta_{12}>0$ and $\operatorname{supp}\mathscr{F}_{d(N-1)}\phi_{2}\subset\{(\xi_{3},\ldots,\xi_{N})\in\mathbb{R}^{d(N-2)}\bigm||\xi_{3}|<1,\ldots,|\xi_{N}|<1\}$. We define $\Phi_{v}=\mathscr{T}_{v}\Phi_{0}$ for
@@
Let $\Phi_{v}$ be defined as above and $\Psi_{v}$ have the same properties for $\Psi_{0}=\psi_{1}\otimes\phi_{2}$. Then
@@
Acknowledgments. This work was supported by JSPS KAKENHI Grant Numbers JP21K03279.

### v4

@@ 1 Introduction
We now give the assumptions for each potential function $V_{jk}$ in (1.1) as a perturbation of $H_{0}(t)$. Let $\lambda$ satisfy
@@
This $\lambda$ describes the effective scaling rate induced by the time-decaying harmonic potential and is directly related to the classical trajectories $x(t)=O(|t|^{1-\lambda})$ as $|t|\rightarrow\infty$. The condition $0<\lambda<1/2$ ensures the short-range character of the interaction under the scaling associated with the propagator. The bracket $\langle\cdot\rangle$ is defined by $\sqrt{1+|\cdot|}$ and $A\lesssim B$ means that there exists a constant $C>0$ such that $A\leqslant CB$.
@@
with $\rho>1/(1-\lambda)$. The singular part $V_{jk}^{\rm sing}\in L^{q}(\mathbb{R}^{d})$ is compactly supported on $\mathscr{X}$, where the Lebesgue exponent $q$ satisfies
@@
acting on $L^{2}(\mathscr{X})$. Here we note $r_{k}-r_{j}\in\mathscr{X}$ for any $1\leqslant j<k\leqslant N$. Under these definitions, the perturbed full Hamiltonian (1.1) is decomposed into
@@
Many researchers have studied scattering theory under the $N$-body quantum systems, in particular, the the existence and asymptotic completeness of the wave operators, has been studied. Regarding the standard $N$-body Schrödinger operators, see [5, 13, 14, 18, 46, 47, 48]; the detailed history and proofs are also provided in [15, 31]. For $N$-body Schrödinger operators in the external electric fields, see [1, 3, 8, 9, 20]. For the external magnetic fields, see [2, 17]. However, there are no results regarding scattering theory for $N$-body time-decaying harmonic oscillators.
@@
Regarding scattering theory for the two-body time-decaying harmonic oscillator, [28] proved the existence of the wave operators under the assumption 1.1 only for the bounded part $V^{\rm bdd}$ and non-existence for $\rho\leqslant 1/(1-\lambda)$, and clarified that the condition $\rho>1/(1-\lambda)$ is short-range. We can intuitively understand the threshold $1/(1-\lambda)$ from the classical motion of the particle $x(t)=c_{1}t^{1-\lambda}+c_{2}t^{\lambda}$ for $t\geqslant T$, which satisfies the Newton equation $({\rm d}^{2}/{\rm d}t^{2})x(t)=-k(t)x(t)$ because $\lambda$ is one of the roots of the quadratic equation $\lambda^{2}-\lambda+\sigma=0$. The critical case $\sigma=1/4$ was studied by [29]. For the inverse square potential, [30] proved the asymptotic completeness of the wave operators. Moreover, [30] proved the Strichartz estimates and applied them to the initial-value problems for the non-linear Schrödinger equations. The non-linear Schrödinger equations and Strichartz estimates with time-decaying harmonic oscillators have also been studied by [33, 34, 35, 36, 37, 38, 39, 40].
In contrast to the time-independent harmonic case or standard free Schrödinger dynamics, the time-decaying quadratic potential induces a nontrivial scaling behavior in the classical trajectories and the quantum propagator. One of the key technical ingredients of this paper is a propagation estimate for the free dynamics associated with the time-dependent harmonic oscillator in the $N$-body framework (Proposition 2.2). Although the estimate itself can be reduced to the two-body case by an appropriate choice of Jacobi coordinates, its formulation in a form suitable for the $N$-body setting is essential for proving the existence of the $N$-cluster wave operators. Once the existence of the wave operators is established, the inverse scattering argument follows the Enss-Weder time-dependent method and reduces the problem to suitable two-body channels. While the inverse problem itself is treated by adapting known high-velocity arguments, the derivation of the propagation estimates (Lemmas 3.2 and 3.3) in the presence of a time-decaying harmonic potential and their extension from the two-body case studied in [26] to the $N$-body setting constitute the main technical novelty of the present work.
@@ 2 Existence of Wave Operators
$\|F(\sqrt{\mu_{jk}}|r_{k}-r_{j}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))e^{-{\rm i}t^{1-2\lambda}D^{2}/(2(1-2\lambda))}\phi\|$
$\lesssim_{\nu}t^{(d/2-\nu)(1-2\lambda)}\|\langle r_{k}-r_{j}\rangle^{\nu}\phi\|$ (2.11)
holds for any $\nu\in\mathbb{N}$, $t\geqslant T$ and $1\leqslant j<k\leqslant N$, where $\mu_{jk}=m_{j}m_{k}/(m_{j}+m_{k})$ and $\lesssim_{\nu}$ means that the constant depends on $\nu$.
@@
where we reconstruct the Jacobi coordinates by relabeling them so that the relative coordinate $y_{1}=r_{k}-r_{j}$ appears as the first Jacobi component for $(j,k)\neq(1,2)$. Because $|\xi_{1}|=|L_{jlk}\xi|\geqslant\epsilon_{0}|\xi|\geqslant\epsilon^{2}$ holds on $\operatorname{supp}\mathscr{F}_{d(N-1)}\phi$, it suffices to prove that
$\|F(|x_{1}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))e^{-{\rm i}t^{1-2\lambda}D_{1}^{2}/(2(1-2\lambda))}\psi\|_{L^{2}(\mathbb{R}^{d})}$
$\lesssim_{\nu}t^{(d/2-\nu)(1-2\lambda)}\|\langle x_{1}\rangle^{\nu}\psi\|_{L^{2}(\mathbb{R}^{d})}$ (2.13)
@@
and integrating by parts, we have
$|F(|x_{1}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))e^{-{\rm i}t^{1-2\lambda}D_{1}^{2}/(2(1-2\lambda))}\psi(x_{1})|$
$\lesssim t^{-\nu(1-2\lambda)}F(|x_{1}|\leqslant\epsilon^{2}t^{1-2\lambda}/(1-2\lambda))\|\langle x\rangle^{\nu}\psi\|_{L^{2}(\mathbb{R}_{x}^{d})}.$ (2.17)
Integrating (2.17) with respect to $x_{1}$ yields (2.13). ∎
@@
$\lesssim t^{(d/2-\nu)(1-2\lambda)}\|\langle r_{k}-r_{j}\rangle^{\nu}\phi\|$ (2.21)
@@
$\chi(x)=\begin{cases}\ 1&\mbox{if}\quad|x|\leqslant 2\\ \ 0&\mbox{if}\quad|x|\geqslant 3.\end{cases}$ (2.22)
@@
Choosing $\epsilon>0$ sufficiently small, we can use Proposition 2.2 again for the first and second terms on the right side of (2.25). By (2.19), (2.20), (2.21), and (2.25). We therefore have
$\|({\rm d}/{\rm d}t)U_{\lambda}(t,T)^{*}e^{-{\rm i}t^{1-2\lambda}D^{2}/(2(1-2\lambda))}\phi\|\lesssim t^{-\rho(1-\lambda)}+t^{(d/2-\nu)(1-2\lambda)}+t^{-2+2\lambda}$ (2.26)
for $t\gg 1$. We note that $\rho(1-\lambda)>1$, $2-2\lambda>1$, and $(\nu-d/2)(1-2\lambda)>1$ for $\nu\gg 1$. By the Cook-Kuroda method ([45, Theorem XI.4]), this completes the proof. ∎
@@ 3 Inverse Scattering
For $v\in\mathbb{R}^{d}$, the normalization is $\hat{v}=v/|v|$. Let $\hat{v},e_{3},\ldots,e_{N}\in\mathbb{R}^{d}$ be unit vectors pointing in mutually different directions. We define $v_{jk}=v_{k}-v_{j}$ for $1\leqslant j<k\leqslant N$ where $v_{1}=-v/2$, $v_{2}=v/2$, and $v_{j}=|v|^{2}e_{j}$ for $3\leqslant j\leqslant N$. Let $\Phi_{0}\in\mathscr{S}(\mathbb{R}^{d(N-1)})$ be $\Phi_{0}=\phi_{1}\otimes\phi_{2}$ such that $\mathscr{F}_{d}\phi_{1}\in C_{0}^{\infty}(\mathbb{R}^{d})$ and $\mathscr{F}_{d(N-2)}\phi_{2}\in C_{0}^{\infty}(\mathbb{R}^{d(N-2)})$ with $\|\phi_{2}\|_{L^{2}(\mathbb{R}^{d(N-2)})}=1$. In addition, we assume that $\operatorname{supp}\mathscr{F}_{d}\phi_{1}\subset\{\xi_{12}\in\mathbb{R}^{d}\bigm||\xi_{12}|<\eta_{12}\}$ for some $\eta_{12}>0$ and $\operatorname{supp}\mathscr{F}_{d(N-2)}\phi_{2}\subset\{(\xi_{3},\ldots,\xi_{N})\in\mathbb{R}^{d(N-2)}\bigm||\xi_{3}|<1,\ldots,|\xi_{N}|<1\}$. We define $\Phi_{v}=\mathscr{T}_{v}\Phi_{0}$ for
@@
Let $\Phi_{v}$ be defined as (3.16) and $\Psi_{v}$ have the same properties for $\Psi_{0}=\psi_{1}\otimes\phi_{2}$. Then
@@
Acknowledgments. This work was supported by JSPS KAKENHI Grant Number JP21K03279.